OpenStudy (anonymous):

Four years ago, Jim's father was 5 times as old as Jim. How old is Jim's father now is Jim is 12? a. 56 b. 44 c. 40 d. 36

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

is it b?

OpenStudy (ranga):

Let x be the age of Jim's father now. Four years ago Jim's father would have been (x - 4) Four years ago Jim would have been 8 (as he is 12 now) Four years ago, Jim's father was 5 times as old as Jim. Therefore, x - 4 = 5 * 8 = 40 Solve for x.

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

PLEASE THIS LAST QUESTION!! I CAN FINISH MY HW.

OpenStudy (ranga):

go ahead post it.

OpenStudy (anonymous):

THANKS :)

OpenStudy (anonymous):

The ratio of 3/4 to 5/2 is a. 10 to 3 b. 15 to 8 c. 3 to 10 d. 8 to 15

OpenStudy (ranga):

Dividing by a fraction is same as multiplying by its reciprocal. (3/4) divided by (5/2) = 3/4 * 2/5 = ?

OpenStudy (anonymous):

its C?

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

THANKS SO MUCH FOR UR HELP!!!!! :)

OpenStudy (ranga):

you are welcome.